# Gyrokinetic flux-driven simulations in mixed TEM/ITG regime using a delta-f PIC scheme with evolving background

### v1

@@ VI Conclusion
To that end, an adaptive background density and temperature scheme was introduced. The background density and temperature was evolved through ad-hoc relaxation equations Eqs.(12)-(14) with associated relaxation rates. Along the simulation, the lower order f.s.a. velocity moments of density and kinetic energy that tend to accumulate in the perturbed distribution function were kept low thanks to continuous transfer to the background Maxwellian. The adaptive scheme was implemented in toroidal geometry using the global gyrokinetic code ORB5 Lanti et al. (2020) to simulate mixed TEM-ITG regime turbulence. The electrons have an upgrade hybrid response Lanti (2019), i.e. when solving the Quasi-Neutrality-Equation (QNE) for the self-consistent electrostatic field the model takes into account the drift-kinetic response of all electrons to the zonal perturbations, while for non-zonal perturbations trapped electrons still respond drift-kinetically, the passing electrons however, adiabatically. Heat source radial profiles for the ions and electrons respectively, derived from previous temperature-gradient-driven runs were used as sources for flux-driven runs presented in this work. The adaptive scheme used a canonical Maxwellian control variate Eq.(LABEL:eq:f0cm), and adapted both density and temperature profiles of ions and electrons independently. A comparison of two methods of calculating the r.h.s. correction to the QNE Eq.(LABEL:eq:qne_corr) was conducted, and it was shown that the correction term is necessary to keep correct zonal flow structures.
@@
When compared to the non-adaptive cases, results of the adaptive cases showed higher heat fluxes and lower zonal flow shearing rate amplitudes. The adaptive cases kept the SNR at quasi-steady values, with greatly reduced standard deviation of marker weights. This is demonstrated to be feasible even with adaptive cases having a quarter of the number of markers used by the non-adaptive (standard) cases (from $256$M to $64$M markers). Only the simulations using the adaptive cases managed to reach quasi-steady state with a single marker loading at initial time. Phase-space volume diagnostics were used to detect phase-space volume depletion at low energies near axis for ions. Though this problem occurred for both non-adaptive and adaptive cases, the latter seemed to nonetheless be less affected. Nonetheless, both schemes still suffer from growing regions in phase space due to marker diffusion that are potentially under-sampled as the simulations run over transport time scales. For scenarios with even greater profile deviation, a resampling of markers is expected to become necessary.
@@
As future work, investigations into the inclusion of parallel flow adaptation could be conducted. Introducing time-dependence into background ion gyrocenter density allows one to de-linearize the ion polarization density term in the QNE. The implications of using such explicitly time-dependent term could be explored. So as to also extend the evolving background approach to electromagnetic simulations, a similar correction term would need to be added to Ampère’s law as the one implemented in the QNE. This would then allow efficient flux-driven simulations of kinetic ballooning, tearing, and internal kink modes. In the presence of fast ions, a time-dependent background could also be useful when simulating Alfvén or energetic particle modes.
@@
Further sophistication to the control variate could be envisaged. A control variate expanded in a set of basis functions could be pursued. Though only used as an offline diagnostic, Ref.Bottino et al. (2022) has expressed in ORB5 the background distribution as an expansion in the space of invariants of the unperturbed trajectories. Representing the background velocity distribution in terms of Hermite and Laguerre polynomials basis functions could be used in tandem with collision operators expressed in such a basis Frei, Ernst, and Ricci (2022). A mixed representation akin to the XGC code Ku, Chang, and Diamond (2009); Ku et al. (2016), where the control variate consists of an analytic function plus a correction term represented on a phase-space grid could be an alternative. Nonetheless, complexity added to background translates to complexity added to the code as a whole. A highly complex control variate risks inheriting the disadvantages of both the particle- and grid-based approaches. It may be best to fall-back to the full-f PIC approach for simulations exhibiting high relative fluctuation amplitudes, i.e. in particular for simulating edge or Scrape-Off-Layer (SOL) conditions.

### v3

@@ VI Conclusion
To that end, an adaptive background density and temperature scheme was introduced. The background density and temperature was evolved through the ad-hoc relaxation equations (12)-(14) with associated relaxation rates. Along the simulation, the lower order f.s.a. velocity moments of density and kinetic energy that tend to accumulate in the perturbed distribution function were kept low thanks to continuous transfer to the background Maxwellian. The adaptive scheme was implemented in toroidal geometry using the global gyrokinetic code ORB5 Lanti et al. (2020) to simulate mixed TEM-ITG regime turbulence. The electrons are modeled with an improved hybrid response Idomura (2016); Lanti et al. (2018); Lanti (2019), i.e. when solving the Quasi-Neutrality-Equation (QNE) for the self-consistent electrostatic field the model takes into account the drift-kinetic of all electrons (trapped and untrapped) to the zonal perturbations, while for non-zonal perturbations trapped electrons still respond drift-kinetically, the passing electrons however, adiabatically. Heat source radial profiles for the ions and electrons, respectively, estimated based on previous temperature-gradient-driven runs were used as local fixed power sources for flux-driven runs presented in this work. The adaptive scheme used a canonical Maxwellian control variate Eq.(LABEL:eq:f0cm), and adapted both density and temperature profiles of ions and electrons independently. When imposing Eq.(LABEL:eq:qnrhs_fsa) the evolution of ion and electron f.s.a. gyrocenter densities are however not independent. A comparison of two methods of calculating the r.h.s. correction to the QNE, Eq.(LABEL:eq:qne_corr), was conducted and it was shown that the correction term is necessary to keep correct zonal flow structures.
@@
When compared to the non-adaptive cases, results of the adaptive cases showed higher heat fluxes and lower zonal flow shearing rate amplitudes. The adaptive cases kept the SNR at quasi-steady values, with greatly reduced standard deviation of marker weights. This is demonstrated to be feasible even with adaptive cases having only a quarter of the number of markers used by the non-adaptive (standard) cases (from $256$M to $64$M markers). Only the simulations using the adaptive cases managed to reach quasi-steady state from a single marker loading at initial time. Phase-space volume diagnostics were used to detect phase-space volume depletion at low energies near the magnetic axis for ions. Though this problem occurred for both non-adaptive and adaptive cases, the latter is significantly less affected. Nonetheless, both schemes still suffer from growing regions in phase space that are potentially under-sampled as the simulations run over transport time scales due to marker diffusion at the velocity sampling boundary. For scenarios with even greater profile deviation, a resampling of markers is expected to become necessary.
@@
As future work, investigations into the benefits of also adapting the background parallel flow could be conducted. Introducing time-dependence into background ion gyrocenter density allows one to partly de-linearize the ion polarization density term in the QNE. The implications of using this time-dependent density in the field equation could be explored. So as to also extend the evolving background approach to electromagnetic simulations, a similar correction term would need to be added to Ampère’s law as the one implemented in the QNE. This would then allow efficient flux-driven simulations of kinetic ballooning, tearing, and internal kink modes. In the presence of fast ions, a time-dependent background could also be useful when simulating in particular Alfvén or energetic particle modes.
@@
Further sophistication to the control variate could be envisaged. A control variate expanded in a set of basis functions could be pursued. Though only used as an offline diagnostic, Ref.Bottino et al. (2022) has expressed in ORB5 the background distribution as a polynomial expansion in the space of invariants of the unperturbed trajectories. Representing the background velocity distribution in terms of Hermite and Laguerre polynomials basis functions could be used in tandem with collision operators expressed in such a basis Frei, Ernst, and Ricci (2022). A mixed representation akin to the XGC code Ku, Chang, and Diamond (2009); Ku et al. (2016), where the control variate consists of an analytic function plus a correction term represented on a phase-space grid could be an alternative. Nonetheless, complexity added to background translates to complexity added to the code as a whole. A highly complex control variate risks inheriting the disadvantages of both the particle- and grid-based approaches. It may be best to fall-back to the full-f PIC approach for simulations exhibiting high relative fluctuation amplitudes, i.e. in particular for simulating edge or Scrape-Off-Layer (SOL) conditions.